OpenStudy (tamara7301):

1. 6( 8 + 9) Which operations are represented in the expression? Evaluate the expression.

OpenStudy (tamara7301):

@Crazy_questions

OpenStudy (anonymous):

23

OpenStudy (tamara7301):

lol

OpenStudy (malcolmmcswain):

Multiplication and addition.

OpenStudy (tamara7301):

ok thx

OpenStudy (anonymous):

who want 2 buy me :P

OpenStudy (tamara7301):

psh not me im good im saving my money to buy candy

OpenStudy (anonymous):

lol

OpenStudy (tamara7301):

wait is he answering

OpenStudy (tamara7301):

@rhettwalker14

OpenStudy (anonymous):

23

OpenStudy (welshfella):

evaluate 6 * 17 = ?

OpenStudy (anonymous):

giv me 3min to get back on the landtop ok

OpenStudy (tamara7301):

102?

OpenStudy (anonymous):

nvm

OpenStudy (welshfella):

right you evaluate the parentheses first = 8+9 = 17 then you multiply by 6

OpenStudy (anonymous):

ABC.........123

OpenStudy (tamara7301):

idk what he was saying

OpenStudy (tamara7301):

yay ur here

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im not good at math

OpenStudy (tamara7301):

wait i gtg do something

OpenStudy (anonymous):

welcome fellow wired-O I'm Talking to every1

OpenStudy (anonymous):

who want a medal?

OpenStudy (anonymous):

me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your weclome

OpenStudy (tamara7301):

i gtg u guys byee

OpenStudy (anonymous):

bye